OpenStudy (anonymous):

ln(x)=1.231 solve for x

OpenStudy (anonymous):

\[e^{ln(x)}=1(x)\]

OpenStudy (anonymous):

write both sides in power of e e^ln(x)=e^1.231 then x=e^1.231

OpenStudy (anonymous):

in other words \[e^{ln}=1\]

OpenStudy (anonymous):

@Tannerbball are you getting this ?